# Do Language Models Understand Time?

### v1

@@ Abstract.
Large language models (LLMs) have revolutionized video-based computer vision applications, including action recognition, anomaly detection, and video summarization. Videos inherently pose unique challenges, combining spatial complexity with temporal dynamics that are absent in static images or textual data. Current approaches to video understanding with LLMs often rely on pretrained video encoders to extract spatiotemporal features and text encoders to capture semantic meaning. These representations are integrated within LLM frameworks, enabling multimodal reasoning across diverse video tasks. However, the critical question persists: Can LLMs truly understand the concept of time, and how effectively can they reason about temporal relationships in videos? This work critically examines the role of LLMs in video processing, with a specific focus on their temporal reasoning capabilities. We identify key limitations in the interaction between LLMs and pretrained encoders, revealing gaps in their ability to model long-term dependencies and abstract temporal concepts such as causality and event progression. Furthermore, we analyze challenges posed by existing video datasets, including biases, lack of temporal annotations, and domain-specific limitations that constrain the temporal understanding of LLMs. To address these gaps, we explore promising future directions, including the co-evolution of LLMs and encoders, the development of enriched datasets with explicit temporal labels, and innovative architectures for integrating spatial, temporal, and semantic reasoning. By addressing these challenges, we aim to advance the temporal comprehension of LLMs, unlocking their full potential in video analysis and beyond.
@@
copyright: none
@@ 1. Introduction
Large language models (LLMs) have brought transformative advancements to artificial intelligence (AI), excelling across a wide array of tasks in natural language processing and computer vision (Bugliarello et al., 2021; Ghosh et al., 2024; Xu and Poo, 2023). Their ability to understand and generate human-like language has enabled groundbreaking applications, from machine translation to image and video captioning (Kim et al., 2024) (see Figure 1, frames from EPIC-KITCHENS-100 (Damen et al., 2020)). More recently, the integration of LLMs into video processing has sparked significant interest, leading to advances in tasks such as action recognition (Wang et al., 2022b, 2024h), anomaly detection (Wang et al., 2019b; Zhang et al., 2024d; Zhu et al., 2024), and video summarization (Islam et al., 2024; Yang et al., 2024a; Liu et al., 2024d; Zohar et al., 2024). However, videos pose unique challenges compared to other modalities due to their dual reliance on both spatial and temporal information (Chen et al., 2024d). Unlike static images, videos capture the dimension of time, embedding sequential dynamics that demand sophisticated reasoning (Vondrick et al., 2016; Liu et al., 2018). Similarly, unlike textual data, videos involve rich, complex visual elements that require intricate modeling (Wang et al., [n. d.]; Chen et al., [n. d.]).
@@
COSMO (Wang et al., 2024i) arXiv 2024 Open-CLIP ViT-L/14(Radford et al., 2021) Text: OPT-IML(Iyer et al., 2022)/RedPajama(Together.xyz, 2023)Mistral(Jiang et al., 2023) Gated cross-attention Contrastive-streamlined multimodal model.
@@
MotionLLM (Chen et al., 2024a) arXiv 2024 LanguageBind(Zhu et al., 2023), VQ-VAE(Zhang et al., 2023b) Text: Vicuna(Chiang et al., 2023) Motion / Video translator Understanding human behaviors from human motions and videos.
@@
Holmes-VAD (Zhang et al., 2024d) arXiv 2024 LanguageBind ViT-L/14(Zhu et al., 2023) Text: LLaMA3-Instruct-70B(AI, 2024) Temporal sampler Multimodal LLM for video anomaly detection.
@@
SliME (Zhang et al., 2024c) arXiv 2024 Adaptive slicing Text: LLaMA3-8B(AI, 2024) MLP & query transformer High-resolution multimodal model.
@@
Oryx (Liu et al., 2024a) arXiv 2024 OryxViT(Zhai et al., 2023) Text: Qwen2-7B/32B(Bai et al., 2023; Yang et al., 2024b) Cross-attention Spatial-temporal model for high-resolution understanding.
Video-XL (Shu et al., 2024) arXiv 2024 CLIP ViT-L(Radford et al., 2021) Text: Qwen2-7B(Bai et al., 2023; Yang et al., 2024b) Visual Summarization Tokens Long-context video understanding model.
@@ 2. Related Work
EPIC-KITCHENS(Damen et al., 2020) 2020 Participant kitchens 432 Video+Text+Audio ~458 Start and end timestamps provided Largest egocentric video dataset
@@ 3. Analysis and Discussion
However, pretrained encoders are not without limitations. First, encoders trained on general datasets may underperform on domain-specific video tasks (Zhu et al., 2024; Chen et al., 2024d). Second, many pretrained encoders prioritize spatial over temporal information, necessitating additional modules, such as temporal transformers, to capture complex temporal dynamics (Chen et al., 2024a; Zhang et al., 2024d). Lastly, large-scale video datasets required for training such encoders are costly to annotate, limiting their ability to encompass diverse or abstract video content.
@@
Table 1 highlights the interaction/fusion mechanisms used in the latest video-LLMs under the ‘Interaction / Fusion mechanism’ column. Figure 3(a) visualizes the distribution of various interaction and fusion techniques, while Figure 3(b) showcases the modalities used in closely related works from 2024 (notably, text modalities are consistently used across these studies). As shown in these plots, the majority of works use projection layers (e.g., linear projection (Li et al., 2024), MLP projection (Zhang et al., 2024a, b; Fu et al., 2024b), semantic translators (Chen et al., 2023b, 2024a)) and cross-attention mechanisms (e.g., cross-modal attention (Yang et al., 2023), gated cross-attention (Wang et al., 2024i)) to facilitate interaction between encoders and LLMs. Temporal-specific mechanisms, such as temporal sampler (Zhang et al., 2024d) and scene-level sequential alignment (Yang et al., 2024a), are also used. Emerging video-LLMs are beginning to explore novel fusion mechanisms, including tracklet-centric approach (Wang et al., 2023a), Perceiver resamplers (Alayrac et al., 2022; Zohar et al., 2024), and MQ-Former (Wang et al., 2024c).
@@ Acknowledgements.
Xi Ding, a Research Assistant with the Temporal Intelligence and Motion Extraction (TIME) Lab at ANU, contributed to this work. TIME Lab is a dynamic research team comprising master’s and honours students focused on advancing video processing and motion analysis. This research was conducted under the supervision of Lei Wang.

### v2

@@ Abstract.
Large language models (LLMs) have revolutionized video-based computer vision applications, including action recognition, anomaly detection, and video summarization. Videos inherently pose unique challenges, combining spatial complexity with temporal dynamics that are absent in static images or textual data. Current approaches to video understanding with LLMs often rely on pretrained video encoders to extract spatiotemporal features and text encoders to capture semantic meaning. These representations are integrated within LLM frameworks, enabling multimodal reasoning across diverse video tasks. However, the critical question persists: Can LLMs truly understand the concept of time, and how effectively can they reason about temporal relationships in videos? This work critically examines the role of LLMs in video processing, with a specific focus on their temporal reasoning capabilities. We identify key limitations in the interaction between LLMs and pretrained encoders, revealing gaps in their ability to model long-term dependencies and abstract temporal concepts such as causality and event progression. Furthermore, we analyze challenges posed by existing video datasets, including biases, lack of temporal annotations, and domain-specific limitations that constrain the temporal understanding of LLMs. To address these gaps, we explore promising future directions, including the co-evolution of LLMs and encoders, the development of enriched datasets with explicit temporal labels, and innovative architectures for integrating spatial, temporal, and semantic reasoning. By addressing these challenges, we aim to advance the temporal comprehension of LLMs, unlocking their full potential in video analysis and beyond. Our paper’s GitHub repository can be found here.
@@
journalyear: 2025copyright: ccconference: Companion Proceedings of the ACM Web Conference 2025; April 28-May 2, 2025; Sydney, NSW, Australiabooktitle: Companion Proceedings of the ACM Web Conference 2025 (WWW Companion ’25), April 28-May 2, 2025, Sydney, NSW, Australiadoi: 10.1145/3701716.3717744isbn: 979-8-4007-1331-6/2025/04ccs: Computing methodologies Motion captureccs: Computing methodologies Neural networksccs: Information systems Language modelsccs: Information systems World Wide Web
@@ 1. Introduction
Large language models (LLMs) have brought transformative advancements to artificial intelligence (AI), excelling across a wide array of tasks in natural language processing and computer vision (Bugliarello et al., 2021; Ghosh et al., 2024; Xu and Poo, 2023). Their ability to understand and generate human-like language has enabled groundbreaking applications, from machine translation to image and video captioning (Kim et al., 2024) (see Figure 1, frames from EPIC-KITCHENS-100 (Damen et al., 2022)). More recently, the integration of LLMs into video processing has sparked significant interest, leading to advances in tasks such as action recognition (Wang et al., 2022b, 2024h), anomaly detection (Wang et al., 2019b; Zhang et al., 2024c; Zhu et al., 2024), and video summarization (Islam et al., 2024; Yang et al., 2024a; Liu et al., 2024d; Zohar et al., 2024). However, videos pose unique challenges compared to other modalities due to their dual reliance on both spatial and temporal information (Chen et al., 2024d). Unlike static images, videos capture the dimension of time, embedding sequential dynamics that demand sophisticated reasoning (Vondrick et al., 2016; Liu et al., 2018). Similarly, unlike textual data, videos involve rich, complex visual elements that require intricate modeling (Wang et al., [n. d.]; Chen et al., [n. d.]).
@@
COSMO (Wang et al., 2024i) arXiv 2024 CLIP ViT-L/14(Radford et al., 2021) Text: OPT-IML(Iyer et al., 2022)/RedPajama(Together.xyz, 2023)Mistral(Jiang et al., 2023) Gated cross-attention Contrastive-streamlined multimodal model.
@@
MotionLLM (Chen et al., 2024a) arXiv 2024 LanguageBind(Zhu et al., 2023), VQ-VAE(Zhang et al., 2023b) Text: Vicuna(Chiang et al., 2023) Modality translator: Motion / Video translator Understanding human behaviors from human motions and videos.
@@
Holmes-VAD (Zhang et al., 2024c) arXiv 2024 LanguageBind ViT-L/14(Zhu et al., 2023) Text: LLaMA3-Instruct-70B(AI, 2024) Temporal sampler Multimodal LLM for video anomaly detection.
@@
Oryx (Liu et al., 2024a) arXiv 2024 OryxViT(Liu et al., 2024a; Zhai et al., 2023) Text: Qwen2-7B/32B(Bai et al., 2023; Yang et al., 2024b) Cross-attention Spatial-temporal model for high-resolution understanding.
Video-XL (Shu et al., 2024) arXiv 2024 CLIP ViT-L(Radford et al., 2021) Text: Qwen2-7B(Bai et al., 2023; Yang et al., 2024b) Visual-language projector Long-context video understanding model.
@@ 2. Related Work
Kinetics-600(Carreira et al., 2018) 2018 YouTube 392,622 Video 10 No Human actions (e.g., sports, tasks)
@@
EPIC-KITCHENS(Damen et al., 2018) 2018 Participant kitchens 432 Video+Text+Audio ~458 Start and end timestamps provided Large-scale egocentric cooking dataset
@@
EPIC-KITCHENS-100(Damen et al., 2022) 2020 Participant kitchens 700 Video+Text+Audio ~514 Start and end timestamps provided Large-scale egocentric cooking dataset
@@ 3. Analysis and Discussion
However, pretrained encoders are not without limitations. First, encoders trained on general datasets may underperform on domain-specific video tasks (Zhu et al., 2024; Chen et al., 2024d). Second, many pretrained encoders prioritize spatial over temporal information, necessitating additional modules, such as temporal transformers, to capture complex temporal dynamics (Chen et al., 2024a; Zhang et al., 2024c). Lastly, large-scale video datasets required for training such encoders are costly to annotate, limiting their ability to encompass diverse or abstract video content.
@@
Table 1 highlights the interaction/fusion mechanisms used in the latest video-LLMs under the ‘Interaction / Fusion mechanism’ column. Figure 3(a) visualizes the distribution of various interaction and fusion techniques, while Figure 3(b) showcases the modalities used in closely related works from 2024 (notably, text modalities are consistently used across these studies). As shown in these plots, the majority of works use projection layers (e.g., linear projection (Li et al., 2024), MLP projection (Zhang et al., 2024a, b; Fu et al., 2024b), semantic translators (Chen et al., 2023b, 2024a)) and cross-attention mechanisms (e.g., cross-modal attention (Yang et al., 2023), gated cross-attention (Wang et al., 2024i)) to facilitate interaction between encoders and LLMs. Temporal-specific mechanisms, such as temporal sampler (Zhang et al., 2024c) and scene-level sequential alignment (Yang et al., 2024a), are also used. Emerging video-LLMs are beginning to explore novel fusion mechanisms, including tracklet-centric approach (Wang et al., 2023a), Perceiver resamplers (Alayrac et al., 2022; Zohar et al., 2024), and MQ-Former (Wang et al., 2024c).
@@ Acknowledgements.
Xi Ding, a Research Assistant with the Temporal Intelligence and Motion Extraction (TIME) Lab at ANU, contributed to this work. This research was conducted under the supervision of Lei Wang.